OpenStudy (anonymous):

y + 3 = 2x y + 4 = 3x If you use the comparison method to solve the given system, what is the resulting equation?

OpenStudy (anonymous):

y + 3 = 2x ==> y = 2x - 3 y + 4 = 3x ==> y = 3x - 4 2x - 3 = 3x - 4